OpenStudy (anonymous):

Suppose a straight 1.60mm -diameter copper wire could just "float" horizontally in air because of the force due to the Earth's magnetic field B⃗ , which is horizontal, perpendicular to the wire, and of magnitude 4.2×10−5T . What current would the wire carry?

OpenStudy (anonymous):

Fb=Fg ILB=mg IkB=2 pi (d/2)^2 g (density copper) I = (pi) (0.0016/2)^2 (9.8) (8.9x10^3) / (4.2x10-5) = 4175.38 A